OpenStudy (calculator):

Let f (x, y, z) = x cos (y z). Find all partial derivatives of the first, second, and third order

OpenStudy (experimentx):

first order dfdx =cos(y*z) dfdx =-x*z*sin(y*z) dfdx =-x*y*sin(y*z) second order dfdx =0 dfdx =-x*z^2*cos(y*z) dfdx =-x*y^2*cos(y*z) third order dfdx =0 dfdx =x*z^3*sin(y*z) dfdx =x*y^3*sin(y*z) Rest:: too much work